# From Freshness to Effectiveness: Goal-Oriented Sampling for Remote Decision Making

### v1

@@ IV-C1 Inner-Layer MDP Given Dinkelbach Parameter $\lambda$
$\tilde{U}_{K+1}(\lambda)$ $=\min_{A_{i},Z_{i}}\bigg{\{}\,g(\gamma^{\text{r}},Z_{i},A_{i};\lambda)$
$\quad\quad\quad+\tau\,\mathbb{E}\big{[}\tilde{V}_{K}(\gamma^{\prime};\lambda)\mid\gamma^{\text{r}},Z_{i},A_{i}\big{]}\bigg{\}},$ (24a)
$\tilde{V}_{K+1}(\gamma;\lambda)$ $=(1-\tau)\,\tilde{V}_{K}(\gamma;\lambda)$
$\quad+\min_{A_{i},Z_{i}}\bigg{\{}\,g(\gamma,Z_{i},A_{i};\lambda)$
$\quad+\tau\,\mathbb{E}\big{[}\tilde{V}_{K}(\gamma^{\prime};\lambda)\mid\gamma,Z_{i},A_{i}\big{]}\bigg{\}}$
$\quad-\tilde{U}_{K+1}(\lambda),\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A},$ (24b)
@@ IV-D Convergence of $\tau$-RVI
which can be solved by the traditional RVI with guaranteed convergence because of its unichain and aperiodic property. Then, comparing (IV-C1) and (32), we can establish the relationship that $\tau\tilde{V}^{\star}(\gamma;\lambda)={V}^{\star}(\gamma;\lambda)$ for $\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$ and $\tilde{U}(\lambda)={U}(\lambda)$. See Appendix LABEL:proof:convergence:theorem2 for the detailed proof. ∎
@@ V-B Primal-Dinkelbach Synchronous Iteration (Proposed OnePDSI)
$\rho_{K+1}$ $=\min_{A_{i},Z_{i}}\Bigg{\{}\frac{q(\gamma^{\text{r}},Z_{i},A_{i})-\kappa\widetilde{W}_{K}(\gamma^{\text{r}})\cdot\mathbb{E}[Y_{i}]}{f(Z_{i})}$
$+\frac{\kappa\mathbb{E}\left[\widetilde{W}_{K}(\gamma^{\prime})\mid\gamma^{\text{r}},Z_{i},A_{i}\right]\cdot\mathbb{E}[Y_{i}]}{f(Z_{i})}\Bigg{\}}+\widetilde{W}_{K}(\gamma^{\text{r}}),$ (39a)
$\widetilde{W}_{K+1}(\gamma)$ $=\min_{A_{i},Z_{i}}\Bigg{\{}\frac{q(\gamma,Z_{i},A_{i})-\kappa\widetilde{W}_{K}(\gamma)\cdot\mathbb{E}[Y_{i}]}{f(Z_{i})}$
$+\frac{\kappa\mathbb{E}\left[\widetilde{W}_{K}(\gamma^{\prime})\mid\gamma,Z_{i},A_{i}\right]\cdot\mathbb{E}[Y_{i}]}{f(Z_{i})}\Bigg{\}}$
$+\widetilde{W}_{K}(\gamma)-\rho_{K+1},\quad\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}.$ (39b)
@@ VI-A Lagrangian Dual Techniques
$H(\lambda;f_{\max})\triangleq$ (44)
$\inf_{\boldsymbol{\psi}}\lim_{\mathrm{n}\rightarrow\infty}\frac{1}{n}{\sum_{i=0}^{n-1}\left\{\mathbb{E}_{\boldsymbol{\psi}}\left[\sum_{t=\it{D}_{i}}^{\it{D}_{i+1}-1}\mathcal{C}(X_{t},A_{i})\right]-\lambda\mathbb{E}_{\boldsymbol{\psi}}\left[Z_{i}+Y_{i+1}\right]\right\}}$
@@
$\mathcal{L}(\boldsymbol{\psi};\theta,\lambda,f_{\max})=\frac{\theta}{f_{\max}}+$ (45)
$\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{i=0}^{n-1}\mathbb{E}_{\boldsymbol{\psi}}\left\{\sum_{t=D_{i}}^{D_{i+1}-1}\mathcal{C}\left(X_{t},A_{i}\right)-(\lambda+\theta)\left(Z_{i}+Y_{i+1}\right)\right\},$
@@
$\phi^{\star}_{\lambda+\theta^{\star}_{\lambda}}=\mathop{\arg\min}_{\phi}\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{i=0}^{n-1}\mathbb{E}_{\phi}\Bigg{\{}\sum_{t=D_{i}}^{D_{i+1}-1}\mathcal{C}\left(X_{t},A_{i}\right)-$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad(\lambda+\theta^{\star}_{\lambda})\left(Z_{i}+Y_{i+1}\right)\Bigg{\}}+\frac{\theta^{\star}_{\lambda}}{f_{\max}},$ (48b)
@@ 6      Solve LP Problem 7 to obtain $Q^{\star}(f_{\max})$;
In the three-layer algorithm, each update of $\lambda$ or $\theta$ necessitates solving the inner-layer MDP using the $\tau$-RVI algorithm. This process incurs a high computational complexity due to the repeated execution of the $\tau$-RVI algorithm required to iteratively search and optimize the parameters $\lambda$ and $\theta$.
@@ VIII-A1 Sampling Policies
    [63, Algorithm 2] presents a low-complexity algorithm to solve $\beta$.

### v2

@@ IV-C1 Inner-Layer MDP Given Dinkelbach Parameter $\lambda$
$\tilde{U}_{K+1}(\lambda)$ $=\min_{A_{i},Z_{i}}\bigg{\{}\,g(\gamma^{\text{r}},Z_{i},A_{i};\lambda)+\tau\,\mathbb{E}\big{[}\tilde{V}_{K}(\gamma^{\prime};\lambda)\mid\gamma^{\text{r}},Z_{i},A_{i}\big{]}\bigg{\}},$ (24a)
$\tilde{V}_{K+1}(\gamma;\lambda)$ $=(1-\tau)\,\tilde{V}_{K}(\gamma;\lambda)+\min_{A_{i},Z_{i}}\bigg{\{}\,g(\gamma,Z_{i},A_{i};\lambda)+\tau\,\mathbb{E}\big{[}\tilde{V}_{K}(\gamma^{\prime};\lambda)\mid\gamma,Z_{i},A_{i}\big{]}\bigg{\}}-\tilde{U}_{K+1}(\lambda),\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A},$ (24b)
@@ IV-D Convergence of $\tau$-RVI
which can be solved by the traditional RVI with guaranteed convergence because of its unichain and aperiodic property. Then, comparing (IV-C1) and (32), we can establish the relationship that $\tau\tilde{V}^{\star}(\gamma;\lambda)={V}^{\star}(\gamma;\lambda)$ for $\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$ and $\tilde{U}(\lambda)={U}(\lambda)$. See Appendix LABEL:proof:convergence:theorem2 for the proof. ∎
@@ V-B Primal-Dinkelbach Synchronous Iteration (Proposed OnePDSI)
$\rho_{K+1}$ $=\min_{A_{i},Z_{i}}\Bigg{\{}\frac{q(\gamma^{\text{r}},Z_{i},A_{i})-\kappa\widetilde{W}_{K}(\gamma^{\text{r}})\cdot\mathbb{E}[Y_{i}]}{f(Z_{i})}+\frac{\kappa\mathbb{E}\left[\widetilde{W}_{K}(\gamma^{\prime})\mid\gamma^{\text{r}},Z_{i},A_{i}\right]\cdot\mathbb{E}[Y_{i}]}{f(Z_{i})}\Bigg{\}}+\widetilde{W}_{K}(\gamma^{\text{r}}),$ (39a)
$\widetilde{W}_{K+1}(\gamma)$ $=\min_{A_{i},Z_{i}}\Bigg{\{}\frac{q(\gamma,Z_{i},A_{i})-\kappa\widetilde{W}_{K}(\gamma)\cdot\mathbb{E}[Y_{i}]}{f(Z_{i})}+\frac{\kappa\mathbb{E}\left[\widetilde{W}_{K}(\gamma^{\prime})\mid\gamma,Z_{i},A_{i}\right]\cdot\mathbb{E}[Y_{i}]}{f(Z_{i})}\Bigg{\}}+\widetilde{W}_{K}(\gamma)-\rho_{K+1},\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}.$ (39b)
@@ VI-A Lagrangian Dual Techniques
$H(\lambda;f_{\max})\triangleq$ $\inf_{\boldsymbol{\psi}}\lim_{\mathrm{n}\rightarrow\infty}\frac{1}{n}{\sum_{i=0}^{n-1}\left\{\mathbb{E}_{\boldsymbol{\psi}}\left[\sum_{t=\it{D}_{i}}^{\it{D}_{i+1}-1}\mathcal{C}(X_{t},A_{i})\right]-\lambda\mathbb{E}_{\boldsymbol{\psi}}\left[Z_{i}+Y_{i+1}\right]\right\}}$ (44)
@@
$\mathcal{L}(\boldsymbol{\psi};\theta,\lambda,f_{\max})=\frac{\theta}{f_{\max}}+\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{i=0}^{n-1}\mathbb{E}_{\boldsymbol{\psi}}\left\{\sum_{t=D_{i}}^{D_{i+1}-1}\mathcal{C}\left(X_{t},A_{i}\right)-(\lambda+\theta)\left(Z_{i}+Y_{i+1}\right)\right\},$ (45)
@@
$\phi^{\star}_{\lambda+\theta^{\star}_{\lambda}}=\mathop{\arg\min}_{\phi}\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{i=0}^{n-1}\mathbb{E}_{\phi}\Bigg{\{}\sum_{t=D_{i}}^{D_{i+1}-1}\mathcal{C}\left(X_{t},A_{i}\right)-(\lambda+\theta^{\star}_{\lambda})\left(Z_{i}+Y_{i+1}\right)\Bigg{\}}+\frac{\theta^{\star}_{\lambda}}{f_{\max}},$ (48b)
@@ 6      Solve LP Problem 7 to obtain $Q^{\star}(f_{\max})$;
In the above three-layer algorithm, each update of $\lambda$ or $\theta$ necessitates solving the inner-layer MDP using the $\tau$-RVI algorithm. This process incurs a high computational complexity due to the repeated execution of the $\tau$-RVI algorithm required to iteratively search and optimize the parameters $\lambda$ and $\theta$.
@@ VIII-A1 Sampling Policies
    [63, Algorithm 2] presents a low-complexity bisection search algorithm to solve $\beta$.